w = (10,-2, 3,-4). the following: a menhen 6 La ③ ut i = (1, 1, -3,2) ut &=(5, 6, 7, 4) Find the components (that is ru su +10 ① Find 11 prosa cull if å = (-4, 3, 2) and ů = 16, 1,4) Find the vector Ñ such that Proja u + W = 10 1 in puhlem 10
Answer #1

Solution [Page 1 we have, U=(111, -3,2) v=(5,6,7,-4) we (10,02,3,-4) ~ - (5,6,7, 04) (1:1, -3,2) 6) sv +3u - 20 = 5(5,6,7,-4)Solution i no (5,2,1), V. (3111-2) ,we (-6, 1,7) 11 urv 11 11 (8,-1,-1) ✓ 8²+ F1 1²+(112 √66 64+1 Al allull alwll (5² + (2110 ua (a). Page 3 . prog au is geves by projau here, an (-4,3,2) 44. (6,114) froja u (6,114) • (-4, 9, 2) ( +4,312) (+4,312 )

Solved using basic properties of Dot products, Projections and Norms.